OpenStudy (anonymous):

olve: negative 1 over 2 x + 2 = -x + 7

sam (.sam.):

Can you draw it out?

sam (.sam.):

\[-\frac{1}{2}x+2=-x+7\] \[x-\frac{1}{2}x=7-2\] \[\frac{1}{2}x=5\] \[x=10\]

OpenStudy (anonymous):

-1/2x+2=-x+7

sam (.sam.):

Yeah x=10